OpenStudy (anonymous):

-(3w+4)+5(w-2)-3(6w-8)=10

OpenStudy (anonymous):

Still struggling with this eh?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Do you want to figure out how to do this on your own?

OpenStudy (anonymous):

Cause the site won't be able to help you much when it comes to taking the test.

OpenStudy (anonymous):

what site? Yes of course I do.

OpenStudy (anonymous):

Im having the worst time with these

OpenStudy (anonymous):

thanx id

OpenStudy (anonymous):

This site.

OpenStudy (anonymous):

I go away for 2 seconds and someone has already spoon fed you the answer. You cannot get better at it this way.

OpenStudy (anonymous):

Polpak I wasent paying attention I need to learn . You know I want to.

OpenStudy (anonymous):

I do. I'm just being annoyed with the tutors here.

OpenStudy (anonymous):

Do you have a min?

OpenStudy (anonymous):

yes. of course I have all time in the world for this.

OpenStudy (anonymous):

Ok, lets go to a whiteboard.. http://www.twiddla.com/573619

OpenStudy (anonymous):

This has happend before nothing is opening up . try again

OpenStudy (anonymous):

you can open a new browser window and paste that address in.

OpenStudy (anonymous):

its not working.

OpenStudy (anonymous):

polpak?

OpenStudy (anonymous):

its not working...

OpenStudy (anonymous):

Hrm..

OpenStudy (anonymous):

http://www.twiddla.com/573634

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

try that 1

OpenStudy (anonymous):

Got it.

OpenStudy (anonymous):

Answer is w=0. You just need to open the expression(distributive property) -3w-4 + 5w -10 - 18w + 24 =10 -16w +10 = 10 -16w = 10-10 -16w=0 w= 0/(-16) = 0